I am having trouble with this equation....3x + 4y=12. I need to find the x and y intercepts plus the third set of values. I need to plot the intercept points on the graph.
Thank You.

Answer by rfer(16322)

Change to y=mx+b form first.
4y=-3x+12
y=-3/4x+3
y int=(0,3)
x int=(4,0)
(8,-3)